OpenStudy (anonymous):

To produce 3.37 liters of a 1.12 M solution of magnesium nitrate, how many grams of magnesium nitrate must be dissolved?

mathslover (mathslover):

I am a little busy, sorry but I think @abb0t and @aaronq may help you out. They are one of the best Chemistry Experts on OpenStudy! Good Luck

OpenStudy (anonymous):

@abb0t ?

OpenStudy (mertsj):

Same thing. Molar mass times molarity times number of liters.

OpenStudy (abb0t):

148.32 \(\sf \frac{g}{mol}\) \(\times\) \(\sf \frac{1.12~mol}{L}\) \(\times ~\sf 3.37~L\) = grams of Mg(NO\(\sf_3\))\( _2\)